# C3VDv2 - Colonoscopy 3D video dataset with enhanced realism

### v1

@@ Abstract
Computer vision techniques have the potential to improve the diagnostic performance of colonoscopy, but the lack of 3D colonoscopy datasets for training and validation hinders their development. This paper introduces C3VDv2, the second version (v2) of the high-definition Colonoscopy 3D Video Dataset, featuring enhanced realism designed to facilitate the quantitative evaluation of 3D colon reconstruction algorithms. 192 video sequences were captured by imaging 60 unique, high-fidelity silicone colon phantom segments. Ground truth depth, surface normals, optical flow, occlusion, six-degree-of-freedom pose, coverage maps, and 3D models are provided for 169 colonoscopy videos. Eight simulated screening colonoscopy videos acquired by a gastroenterologist are provided with ground truth poses. The dataset includes 15 videos featuring colon deformations for qualitative assessment. C3VDv2 emulates diverse and challenging scenarios for 3D reconstruction algorithms, including fecal debris, mucous pools, blood, debris obscuring the colonoscope lens, en-face views, and fast camera motion. The enhanced realism of C3VDv2 will allow for more robust and representative development and evaluation of 3D reconstruction algorithms.
@@
Colorectal cancer (CRC) is a significant public health concern, ranking as the second leading cause of cancer-related deaths in the United States [1]. Colonoscopy is the gold standard procedure for identifying and resecting precancerous lesions in the gastrointestinal tract [2]. The timely detection and removal of precancerous lesions through routine colonoscopy screening is estimated to reduce the CRC mortality rate by 53%[3]. The protective value of screening colonoscopy depends on the operator’s ability to thoroughly inspect the full colon surface and identify all clinically relevant lesions. A meta-analysis of over 15,000 tandem colonoscopies revealed that approximately 26% of adenomas go undetected [4]. In a retrospective analysis, McGill et al. [5] estimated that 19% of the colon surface area typically remains unobserved during a colonoscopy, potentially leading to missed lesions.
@@
3D computer vision algorithms, including Simultaneous Localization and Mapping (SLAM), Neural Radiance Fields (NeRFs), and 3D Gaussian Splatting, have been used to reconstruct the colon surface from colonoscopy videos [6, 7, 8, 9]. Due to the enclosed nature of the colon, unobserved regions appear as ‘gaps’ in 3D reconstructions, which could highlight areas missed by clinicians [10]. Real-time systems employing these techniques could assist navigation during colonoscopies and increase mucosal coverage. Furthermore, advanced quality metrics, like fractional coverage area, can better assess screening effectiveness. SLAM algorithms include submodules such as camera tracking [11, 12] and depth estimation [13, 14]. These submodules also enable downstream applications such as anatomical landmark recognition [15], using depth to improve lesion classification [16], polyp size estimation [17], and augmented reality for training novice gastroenterologists [18].
@@
3D reconstruction of the colon from colonoscopy video data presents several important challenges for computer vision algorithms. These include the absence of distinct and repetitive features due to the textureless appearance of the colon surface, specular reflections, soft tissue deformation from peristalsis or distension, and occlusions caused by haustral folds [19]. Additional difficulties arise from the presence of fecal debris, rapid scope motion, limited viewpoints within the constrained tubular anatomy, wide range of working distances present in many scenes, and non-uniform illumination from the co-located light source and camera. Despite their clinical relevance, such challenging factors are rarely captured together in existing datasets that provide quantitative 3D ground truth. Thus, there is a need for datasets providing pixel-wise 3D ground truth for realistic and challenging colonoscopy scenarios.
@@
There are limited options for obtaining in vivo colon 3D ground truth information. Current monocular colonoscopy datasets derive 3D information from Computed Tomography (CT) scans of the human colon [20, 21, 22, 23], and optical scanning of ex vivo animal tissue [24]. Each approach offers a trade-off between realism and ground truth accuracy. Although datasets imaging ex vivo [24] or in vivo tissue [25] provide realistic videos, they lack pixel-wise 3D ground truth. Conversely, simulated datasets generated with rendering engines offer accurate pixel-level ground truth but struggle to replicate real-world camera optics, illumination models, and light-tissue interactions [23, 26, 27, 20, 24, 21]. Moreover, most synthetic datasets fail to incorporate realistic conditions such as fecal debris, blood, colon deformations, and the presence of endoscopy instruments. A recent synthetic dataset has made progress in this area by providing realistic vasculature, polyps, ulcers, and blood simulations, but it primarily focuses on lesion segmentation and classification, and thus lacks 3D ground truth [28].
@@
To address these limitations, Bobrow et al. [29] introduced the Colonoscopy 3D Video Dataset (C3VD), which utilizes high-fidelity silicone phantoms with known 3D shapes, imaged with a high-definition clinical colonoscope. C3VD’s primary advantage over real colonoscope datasets, such as EndoSLAM [24] and EndoMapper [25], is providing per-frame pixel-level registered ground truth data. By imaging phantoms with a clinical colonoscope, C3VD inherently captures non-global illumination, real fisheye intrinsics, and non-linear post-processing effects, thereby addressing key limitations of simulated datasets [23, 20, 24, 21]. However, the original C3VD videos only spanned idealized scenarios with clean colons and lenses, slow and smooth camera motion, a static colon, and no tools and accessories present in the field of view (FoV).
C3VDv2 retains the advantages of C3VD, including clinical HD colonoscope videos with registered ground pixel-wise depth maps, normals, optical flow, occlusion maps, diffuse images, and camera poses. Additionally, it introduces several enhancements over the original C3VD dataset to improve realism and comprehensiveness, and increases the dataset size. More specifically, the C3VDv2 dataset incorporates the following improvements:
@@
    Larger dataset: The number of videos has increased eightfold (n=192). The variety of colon geometries has expanded twofold, with eight full colon phantoms.
@@
    Realistic artifacts: The realism of colon appearance has been significantly improved by including artifacts such as fecal debris, mucus pools, blood, foam, and debris and water on the lens. Furthermore, C3VDv2 incorporates instruments like water jets, lens cleaning, and suction.
@@
    Challenging scenarios: The dataset also includes scenarios that are particularly challenging for 3D reconstruction and tracking algorithms. These include fast and abrupt camera motion, en-face to down-the-barrel pose transitions, close-up en-face views of textureless surfaces, the endoscope getting covered in debris, and lens cleaning. The dataset also includes trajectories such as straight-line in-and-out motions, loops where the first and last points are the same, and paths where the first half mirrors the second half with lens cleaning in the middle.
@@
    Paired clean & debris-filled colon frames: Lastly, for every debris-filled colon video, there is a corresponding clean colon video acquired by imaging a colon phantom with no debris using an identical camera trajectory as the debris-filled video. These paired videos enable the quantification of performance degradation due to debris.
C3VDv2 is a valuable resource for training and validating endoscopy 3D reconstruction algorithms. The dataset is publicly available and includes all 3D models, camera pose trajectories, coverage mesh, and occlusion maps, essential for validating SLAM algorithms to identify missing regions. Pixel-level depth estimation, surface normals, and optical flow data can be used to validate both deep learning algorithms and components of SLAM methods. Relative to the first version of C3VD, C3VDv2 dramatically expands the diversity of realistic scenes while increasing the dataset size, enabling comprehensive quantitative evaluation of computer vision algorithms across a wide range of realistic and challenging colonoscopy scenarios.
@@ Methods
The C3VDv2 dataset includes colonoscopy videos with pixel-level ground truth depth maps. To achieve this, we first established a protocol for fabricating high-fidelity phantoms from a 3D surface model of a human colon. Then, we record video sequences of a colonoscope advanced through colon phantoms with a robotic arm while recording the pose log. Finally, we used a multimodal 2D/3D video registration technique to align the acquired video and trajectory data with a ground truth 3D surface model of the colon. A more detailed description of the registration method is described in C3VD [29].
@@
To develop realistic colon phantoms for colonoscopy research, two complete 3D colon models were sculpted using MeshMixer software based on colonoscopy videos and the C3VD model. Each model was reviewed by an experienced gastroenterologist for anatomical accuracy. Colon 1 (c1) was divided into eight distinct regions, while Colon 2 (c2) comprised seven regions, aligning with anatomical sections - cecum, ascending colon, transverse colon, descending colon, sigmoid colon, and rectum; further subdivisions such as transverse 1 and transverse 2 were created for longer segments to fit the 3D printer’s build volume (14.5 × 14.5 × 19.3 cm). Lesions representing various categories of the Paris Classification [30] were sculpted onto the models and strategically placed such that the haustral folds would occlude some polyps. Comprehensive details regarding the lengths of model segments, types of lesions, and their major axis diameters are presented in Tables 2 and 3, and an illustration is presented in Figure 3.
@@
Silicone (Dragon Skin, Smooth-On, Inc.) was used to create casts from each mold, with color and texture variations introduced using silicone pigments (Silc Pig, Smooth-On, Inc.). For each phantom, the initial layer of silicone was manually layered onto the insert. 5 to 12 layers were applied to allow for detailed hand-painting of features such as vasculature patterns on intermediate layers. Silicone was later poured into the bulk of the mold through inlets in the outer molds, filling any gaps between the insert and outer shells. This entire process is illustrated in Figure 2.
Using the molds for each model segment, we cast four complete phantoms, each varying in texture and color. This effort resulted in a total of eight full-length colon phantoms as shown in Figure 4. The diversity in tissue colors and textures among these phantoms was explicitly designed to enable domain randomization for robust training and evaluation of computer vision algorithms.
@@
The experimental setup employed for data acquisition is illustrated in Figure 2. We used an Olympus CF-HQ190L video colonoscope, CV-190 video processor, and CLV-190 light source to image the phantom models. Uncompressed video frames were captured from the clinical video processor using an Orion HD (Matrox Imaging) graphics adapter. Video frames were parsed and stored in raw uncompressed RGB format. Silicone models were placed inside the outer shell molds to prevent phantom motion and deformation during video recording. All videos were recorded using white light illumination. The colonoscope’s distal tip was rigidly and securely mounted to the end effector of a Universal Robotics UR-3e robotic arm. 6-12 waypoints were manually programmed for each video segment to simulate colonoscopy trajectories. The robotic arm traversed the colon by interpolating between these poses, with a repeatability of ±$\pm 0.03$ mm. A pose log was recorded from the arm at a sampling rate of 500 Hz.
@@
Colonoscopy artifacts, such as specular reflections, mucous pools, fecal debris, blood, and foam, pose significant challenges for 3D reconstruction algorithms. We introduce these artifacts to increase the realism of the colonoscopy videos. Real colonoscopy videos served as a reference for creating the debris-filled colon environment. To emulate the specular reflections of the mucosa, the phantom was coated with a silicone lubricant (015594011516, BioFilm, Inc.). A mixture of peanut butter (051500255377, Jif), applesauce (085239045145, Good & Gather), and yellow pigment powder (B0B82NLDDR, Rock N Soil) mimicked fecal debris. Different grades of bowel preparation were created by varying the quantities of debris. Simulated blood was created by combining glycerine (754207111907, Raw Plus Rare), red dye (B089HYVC3R, Craft County), and applesauce. Mucous pools were simulated using a mixture of glycerine, yellow dye, and chicken soup (051000134592, Campbell). Soap foam, colored with yellow pigment powder or red dye, was also added to the debris. The colonoscope’s auxiliary water jet system was used to flush debris from the phantom’s walls. Water-on-lens and lens cleaning scenarios were captured using the colonoscope’s lens cleaning system. In simulated screening videos, the system’s suction feature was used to drain mucous pools and debris. Representative examples are shown in Figure 1 and 5. These debris formulations were reviewed by an experienced gastroenterologist for qualitative agreement with clinical scenes.
Pixel-level GT Video Collection Protocol
@@
In addition to the short videos featuring phantom segments, we created simulated screening colonoscopy videos of complete colon phantoms. Individual phantom segments were bonded together using a silicone adhesive (Sil-Poxy™, Smooth-On, Inc.) and selectively sutured for structural reinforcement. As discussed previously, materials such as glycerin and yellow dye were incorporated to simulate debris. The full colon phantom model was mounted on a cutout foam scaffold to reduce global movement and colon deformation during imaging.
For camera tracking in this experiment, a six-degree-of-freedom electromagnetic (EM) sensor (Aurora, Northern Digital Inc.) was affixed to the distal end of the colonoscope, and pose was sampled at 40 Hz. An experienced gastroenterologist (VA) performed the simulated screening colonoscopy to capture realistic camera trajectories. In total, eight full-colon simulated screening colonoscopy videos (withdrawal phase) were recorded, each corresponding to different texture phantoms and debris. The average duration for these videos was approximately 6.5 minutes, consistent with the >6-minute withdrawal phase duration guideline issued by the American College of Gastroenterology taskforce [31].
2D-3D Registration Pipeline
The 2D-3D registration process proposed in C3VD [29] was followed. A virtual camera was traversed along the recorded real colonoscope trajectory to generate pixel-level ground truth for each video frame. This virtual camera rendered corresponding ground truth frames from the 3D model of the colon. A key challenge is determining the relative location of the phantom model in relation to the virtual camera’s trajectory. Assuming the phantom model remained stationary during the video capture, its pose can be represented by a single, rigid body transformation. To estimate this unknown transformation, we utilize a 2D-3D registration approach. This approach aligns the virtual 3D model with the 2D video frames using extracted edge maps. The registration process iteratively samples the model transform parameter space and measures alignment between edge maps from the video frames and depth maps rendered by the virtual camera.
@@
The recorded video frames and pose log are preprocessed before the registration step. Five keyframes were uniformly sampled from each video sequence. This number was empirically chosen to balance registration accuracy and computational efficiency in C3VD [29]. When the selected keyframes lacked informative edges (e.g., close-up of textureless regions) or contained distorted edges (e.g., lens cleaning artifacts), the closest temporally adjacent frame exhibiting clear edges was manually selected as a replacement. After selecting keyframes, preprocessing was applied to facilitate the generation of high-quality edge maps. First, specular reflections were removed using Telea inpainting [32] to reduce spurious edges. Due to the diverse range of tissue colors, textures and lighting conditions present in the phantom videos, contrast enhancement was necessary. Key frame images were first denoised with a Gaussian filter, followed by grayscale conversion. Contrast Limited Adaptive Histogram Equalization (CLAHE) [33] was then applied. Subsequently, the deep learning-based edge extraction method DexiNed[34] was utilized. The pre-trained model weights (BIPED_10.pt) available from the DexiNed repository were used for this purpose. As the DexiNed output edge images utilized an inverted format (white background, black edges), key frame edges were first inverted. Finally, non-maximal suppression and hysteresis thresholding submodules from the Canny edge detection[35] method were applied. These preprocessing steps generated binary edge maps with thin edges from the input RGB key frames.
@@
The pose log underwent preprocessing to reduce noise. A box filter with a window size of 54 ms was applied to the data. Subsequently, the pose log was downsampled to 60 Hz. It is important to note that the pose log records the position of the robotic arm’s end-effector relative to its base, not the camera itself. Hence, a hand-eye calibration was performed to estimate the camera poses. This calibration established a transformation matrix $\mathbf{X}$ that related the relative motion of the robotic arm $\mathbf{A}$ to the relative motion of the colonoscope camera $\mathbf{B}$. This relationship was expressed by the equation $\mathbf{A}\mathbf{X}=\mathbf{X}\mathbf{B}$. The Park & Martin [36] optimization method was used to solve for the unknown transformation $\mathbf{X}$. To synchronize the video frames and the pose log, the time delay between the start of motion in the video and the pose was calculated. Synchronization was achieved by determining the relative temporal offset that maximized the correlation between the magnitude of optical flow (indicating camera motion) and the displacement recorded in the translation components of the pose. Finally, after transforming robotic arm poses into camera poses and synchronizing them with the video frames, the camera pose for each keyframe was computed by interpolation between the temporally closest camera poses.
@@
For the simulated screening videos, temporal synchronization between video frames and EM tracker poses was established by manually identifying correspondence points, training a linear regression model (time_sensor = drift x time_video + offset) to correct clock drift, then interpolating poses to align with adjusted video frame timestamps. Clock drift arises due to the lack of hardware synchronization between the colonoscope’s video acquisition and the EM tracker, which operate on separate internal clocks. Over time, small differences in their clock frequencies accumulate, leading to misalignment between recorded timestamps. The linear model compensates for this temporal drift, enabling accurate alignment of pose data with corresponding video frames. Then, a hand-eye transform was applied to derive the camera poses from the EM tracker poses, using the same method described above.
@@
An essential feature of clinical colonoscopes is their wide FoV, which is typically around $170^{\circ}$. This feature maximizes the colon surface area visible during a colonoscopy procedure. C3VD incorporated this fisheye effect in its virtual camera model to render the pixel-wise ground truth. As in C3VD, we use the spherical omnidirectional camera model proposed in Scaramuzza et al. [37].
The spherical camera model establishes a one-to-one correspondence between a pixel $(u",v")$ on the image plane and a ray direction in 3D world space $(X,Y,Z)$. The Scaramuzza model employs a polynomial projection function to map the ideal undistorted image coordinates $(u,v)$ to the ray direction, by:
@@
where $\lambda$ is a scalar factor, $\alpha_{i}$ are the polynomial coefficients ($\alpha_{1}=0$), and $\rho=\sqrt{u^{2}+v^{2}}$ is the radial distance of the pixel from the image center. To account for lens-sensor misalignment and distortion, the ideal image coordinates are transformed into real distorted image pixels $(u",v")$ using the following equation:

### v2

@@ Abstract
Spatial computer vision techniques have the potential to improve the diagnostic performance of colonoscopy. However, the lack of 3D colonoscopy datasets for training and validation hinders their development. This paper introduces C3VDv2, the second version (v2) of the high-definition Colonoscopy 3D Video Dataset, featuring enhanced realism designed to facilitate the quantitative evaluation of 3D colon reconstruction algorithms. 192 video sequences totaling 169,371 frames were captured by imaging 60 unique, high-fidelity silicone colon phantom segments. Ground truth depth, surface normals, optical flow, occlusion, diffuse maps, six-degree-of-freedom pose, coverage map, and 3D models are provided for 169 colonoscopy videos. Eight simulated screening colonoscopy videos acquired by a gastroenterologist are provided with ground truth poses. Lastly, the dataset includes 15 videos with colon deformations for qualitative assessment. C3VDv2 emulates diverse and challenging scenarios for 3D reconstruction algorithms, including fecal debris, mucous pools, blood, debris obscuring the colonoscope lens, en-face views, and fast camera motion. The enhanced realism of C3VDv2 will allow for more robust and representative development and evaluation of 3D reconstruction algorithms.
@@
Colorectal cancer (CRC) is a significant public health concern, ranking as the second leading cause of cancer-related deaths in the United States [1]. Colonoscopy is the gold standard procedure for screening CRC in the gastrointestinal tract [2]. The timely detection and removal of precancerous lesions through routine colonoscopy screening is estimated to reduce the CRC mortality rate by 53% [3]. The protective value of screening colonoscopy depends on the operator’s ability to thoroughly inspect the full colon surface and identify all clinically relevant lesions. A meta-analysis of over 15,000 tandem colonoscopies revealed that approximately 26% of adenomas go undetected [4]. In a retrospective analysis, McGill et al. [5] estimated that 19% of the colon surface area typically remains unobserved during a colonoscopy, offering an explanation for a fraction of these missed lesions.
@@
3D computer vision algorithms, including Simultaneous Localization and Mapping (SLAM), have been used to reconstruct the colon surface from colonoscopy videos [6, 7]. Because of the enclosed, tubular anatomy of the colon, missed regions appear as ‘gaps’ in 3D reconstructions, and these gaps can be flagged to help clinicians provide more complete examinations [8]. Real-time systems employing these techniques could assist navigation during colonoscopies and increase mucosal coverage. Furthermore, advanced quality metrics derived from reconstructed topography, like fractional coverage area, can better assess screening effectiveness.
@@
3D reconstruction of the colon from colonoscopy video data presents several important challenges for computer vision algorithms. These include the lack of distinct features on the smooth colon surface, specular reflections, soft tissue deformation from peristalsis or distension, and occlusions caused by haustral folds [9]. Additional difficulties arise from the presence of fecal debris, rapid scope motion, limited viewpoints within the constrained tubular anatomy, a wide range of working distances present in many scenes, and non-uniform illumination from the co-located light source and camera. Despite their clinical relevance, such challenging factors are rarely captured together in existing datasets that provide quantitative 3D ground truth (GT). Thus, there is a need for datasets providing quantitative 3D ground truth for realistic and challenging colonoscopy scenarios.
@@
There are limited options for obtaining in vivo colon 3D ground truth information. Current monocular colonoscopy datasets derive 3D information from Computed Tomography (CT) scans of the human colon [10, 11, 12, 13], and optical scanning of ex vivo animal tissue [14]. Each approach offers a trade-off between realism and ground truth accuracy. Although datasets imaging ex vivo [14] or in vivo tissue [15] provide realistic videos, they lack pixel-wise 3D ground truth. Conversely, simulated datasets generated with rendering engines offer accurate pixel-level ground truth but struggle to replicate real-world camera optics, illumination models, and light-tissue interactions [13, 16, 17, 10, 12, 14, 11]. Moreover, most synthetic datasets fail to incorporate realistic conditions such as fecal debris, blood, colon deformations, and the presence of endoscopy instruments. A recent synthetic dataset has made progress in this area by providing realistic vasculature, polyps, ulcers, and blood simulations, but it primarily focuses on lesion segmentation and classification, and thus lacks 3D ground truth [18]. A summary of existing monocular colonoscopy datasets used in computer vision is presented in Table 1.
@@
To address these limitations, Bobrow et al. [19] introduced the Colonoscopy 3D Video Dataset (C3VD), which utilizes high-fidelity silicone phantoms with known 3D shapes, imaged with a high-definition clinical colonoscope. Unlike real colonoscope datasets, such as EndoSLAM [14] and EndoMapper [15], C3VD provides per-frame pixel-level registered ground truth data while still capturing realistic illumination, fisheye optics, and post-processing artifacts that simulated datasets lack [13, 10, 14, 11]. However, the original C3VD videos only spanned idealized scenarios with clean colons and lenses, slow and smooth camera motion, a static colon, and no tools and accessories present in the field of view (FoV).
C3VDv2 retains all the advantages of the C3VD while introducing several key enhancements to improve realism, comprehensiveness, and scale:
@@
    Larger dataset: The number of videos has increased sevenfold (n=192). The variety of colon geometries has doubled, with eight full colon phantoms.
@@
    Realistic artifacts: The realism of colon appearance has been significantly improved by including artifacts such as fecal debris, mucus pools, blood, foam, and debris and water on the lens (figure 1). Furthermore, C3VDv2 incorporates instruments like water jets, lens cleaning, and suction.
@@
    Challenging scenarios: The dataset includes scenarios that are particularly challenging for 3D reconstruction and tracking algorithms. These include fast and abrupt camera motion, en-face to down-the-barrel pose transitions, close-up en-face views of textureless surfaces, lens occlusion from debris, and lens cleaning. The dataset also includes trajectories such as straight-line in-and-out motions, loops where the first and last points are the same, and mirrored paths with lens cleaning midway (figure 2).
@@
    Paired clean & debris-filled colon frames: For every debris-filled colon video, there is a corresponding clean colon video acquired by imaging a colon phantom with no debris using an identical camera trajectory as the debris-filled video. These paired videos enable the quantification of performance degradation due to debris.
C3VDv2 is a valuable resource for the development and validation of a wide range of 3D computer vision algorithms. For example, it provides the ground truth 3D models, camera poses, and coverage meshes needed to quantitatively validate endoscopy SLAM algorithms [20, 21, 22]. The availability of camera poses also makes C3VDv2 suitable for volumetric rendering techniques such as Neural Radiance Fields (NeRFs) [23, 24] and 3D Gaussian Splatting [25, 26]. Furthermore, the pixel-level depth, surface normal, and optical flow data support the development of SLAM sub-components like depth estimation [27, 28, 29, 30] and camera tracking [31, 11]. These tasks enable downstream applications like anatomical landmark recognition [32], improved lesion classification [33], polyp size estimation [34, 35], and augmented reality for training residents [36]. The dataset also includes simulated screening videos performed by an experienced gastroenterologist on full colon phantoms, providing realistic camera trajectories for validation of camera pose estimation methods. The diversity in tissue color, texture, and camera trajectories enables testing for domain generalization and robustness of these algorithms.
C3VD’s high-resolution 3D models and phantom-making protocols are open-sourced, enabling researchers to modify, 3D print colon molds, and cast their own phantoms for imaging. These protocols can be expanded to fabricate other tissue phantoms, such as other regions of the gastrointestinal tract, like the esophagus and stomach, or to create organ structures for surgical datasets. Beyond 3D reconstruction, C3VDv2 can also be used for auxiliary colonoscopy tasks, such as artifact detection [37] and haustral fold detection [38, 39] in the presence of debris. The dataset could also be used to generate synthetic data [40, 41] for semi- or self-supervised learning [42]. Researchers may further annotate bowel-preparation quality of videos using the Boston scale [43] to gauge performance across varying levels of debris. Since text labels describing artifacts and events in the videos are also provided, the video-text data could aid training medical vision-language models [44]. C3VDv2 also holds potential for polyp detection and tracking. Some polyps have been strategically positioned to be occluded by haustral folds. Lastly, fecal debris could generate false-positive detections, presenting a challenging validation set for lesion detection algorithms.
In summary, C3VDv2 substantially extends the scale and realism of the original dataset, enabling comprehensive evaluation of computer vision methods under diverse and clinically relevant colonoscopy conditions.
@@ Methods
The C3VDv2 dataset includes colonoscopy videos with pixel-level ground truth maps. To achieve this, we first established a protocol for fabricating high-fidelity phantoms from a 3D surface model of a human colon. Then, we record video sequences of a colonoscope advanced through colon phantoms with a robotic arm while recording the pose log. Finally, we used a multimodal 2D/3D video registration technique to align the acquired video and trajectory data with a ground truth 3D surface model of the colon. A more detailed description of the registration method is described in C3VD [19].
@@
To develop realistic colon phantoms for colonoscopy research, two complete 3D colon models were sculpted using MeshMixer software (AutoDesk, Inc.) based on colonoscopy videos and the C3VD model. Each model was reviewed by an experienced gastroenterologist (V.S.A.) for anatomical accuracy. Colon 1 (c1) was divided into eight distinct regions, while Colon 2 (c2) comprised seven regions, aligning with anatomical sections - cecum, ascending colon, transverse colon, descending colon, sigmoid colon, and rectum; further subdivisions such as transverse 1 and transverse 2 were created for longer segments to fit the 3D printer’s build volume (14.5 × 14.5 × 18.5 cm). Lesions representing various categories of the Paris Classification [45] were sculpted onto the models and strategically placed such that the haustral folds would occlude some polyps. Comprehensive details regarding the lengths of model segments, types of lesions, and their major axis diameters are presented in Tables 2 and 3, and an illustration is presented in figure 4.
@@
Silicone (Dragon Skin, Smooth-On, Inc.) was used to create casts from each mold, with color and texture variations introduced using silicone pigments (Silc Pig, Smooth-On, Inc.). For each phantom, the initial layers of silicone were manually painted onto the insert. 5 to 12 layers were applied to allow for detailed hand-painting of features such as vasculature patterns on intermediate layers. Silicone was later poured into the bulk of the mold through inlets in the outer molds, filling any gaps between the insert and outer shells. This entire process is illustrated in figure 3.
Using the molds for each model segment, four complete phantoms were cast, each varying in texture and color. This effort resulted in a total of eight full-length colon phantoms as shown in figure 5. The diversity in tissue colors and textures among these phantoms was explicitly designed to enable domain randomization for robust training and evaluation of computer vision algorithms.
Future versions of these phantoms could incorporate tuned optical scattering and absorption properties to simulate more realistic light–tissue interactions [46, 47, 48]. Additionally, phantoms embedded with polyps suitable for simulated polypectomy procedures [49] would support the integration of therapeutic endoscopy tools in the dataset.
@@
The experimental setup employed for data acquisition is illustrated in figure 3. An Olympus CF-HQ190L video colonoscope, CV-190 video processor, and CLV-190 light source were used to image the phantom models. Uncompressed video frames were captured from the clinical video processor using an Orion HD (Matrox Imaging) graphics adapter. Video frames were parsed and stored in raw uncompressed RGB format. Silicone models were placed inside the outer shell molds to prevent phantom motion and deformation during video recording. All videos were recorded using white light illumination. The colonoscope’s distal tip was rigidly and securely mounted to the end effector of a Universal Robotics UR-3e robotic arm. 6-12 waypoints were manually programmed for each video segment to simulate colonoscopy trajectories. The robotic arm traversed the colon by interpolating between these poses, with a repeatability of $\pm 0.03$ mm. A pose log was recorded from the arm at a sampling rate of 500 Hz.
@@
Colonoscopy artifacts, such as specular reflections, mucous pools, fecal debris, blood, and foam, pose significant challenges for 3D reconstruction algorithms. These artifacts were introduced to increase the realism of the colonoscopy videos. Real colonoscopy videos served as a reference for creating the debris-filled colon environment. To emulate the specular reflections of the mucosa, the phantom was coated with a silicone lubricant (015594011516, BioFilm, Inc.). A mixture of peanut butter (051500255377, Jif), applesauce (085239045145, Good & Gather), and yellow pigment powder (B0B82NLDDR, Rock N Soil) mimicked fecal debris. Different grades of bowel preparation were created by varying the quantities of debris. Simulated blood was created by combining glycerine (754207111907, Raw Plus Rare), red dye (B089HYVC3R, Craft County), and applesauce. Mucous pools were simulated using a mixture of glycerine, yellow dye, and chicken soup (051000134592, Campbell). Soap foam, colored with yellow pigment powder or red dye, was also added to the debris. The colonoscope’s auxiliary water jet system was used to flush debris from the phantom’s walls. Water-on-lens and lens cleaning scenarios were captured using the colonoscope’s lens cleaning system. In simulated screening videos, the system’s suction feature was used to drain mucous pools and debris. Representative examples are shown in figure 1 and 6. These debris formulations were reviewed by an experienced gastroenterologist for qualitative agreement with clinical scenes.
Pixel-level GT & Deformation Video Collection Protocol
@@
In addition to the short videos featuring phantom segments, simulated screening colonoscopy videos of complete colon phantoms were recorded. Individual phantom segments were bonded together using a silicone adhesive (Sil-Poxy™, Smooth-On, Inc.) and selectively sutured for structural reinforcement. As discussed previously, materials such as glycerin and yellow dye were incorporated to simulate debris. The full colon phantom model was mounted on a cutout foam scaffold to reduce global movement and colon deformation during imaging.
For camera tracking in this experiment, a six-degree-of-freedom electromagnetic (EM) sensor (Aurora, Northern Digital Inc.) was affixed to the distal end of the colonoscope, and poses were sampled at 40 Hz. An experienced gastroenterologist (V.S.A.) performed the simulated screening colonoscopy to capture realistic camera trajectories. In total, eight full-colon simulated screening colonoscopy videos (withdrawal phase) were recorded, each corresponding to different texture phantoms and debris. The average duration for these videos was approximately 6.5 minutes, consistent with the >6-minute withdrawal phase duration guideline issued by the American College of Gastroenterology taskforce [50].
2D/3D Registration Pipeline
The 2D/3D registration process proposed in C3VD [19] was followed. A virtual camera was traversed along the recorded real colonoscope trajectory to generate pixel-level ground truth for each video frame. This virtual camera rendered corresponding ground truth frames from the 3D model of the colon. A key challenge is determining the relative location of the phantom model in relation to the virtual camera’s trajectory. Assuming the phantom model remained stationary during the video capture, its pose can be represented by a single, rigid body transformation. To estimate this unknown transformation, a 2D/3D registration approach was utilized. This approach aligns the virtual 3D model with the 2D video frames using extracted edge maps. The registration process iteratively samples the model transform parameter space and measures alignment between edge maps from the video frames and depth maps rendered by the virtual camera.
@@
The recorded video frames and pose log are preprocessed before the registration step. Five keyframes were uniformly sampled from each video sequence. This number was empirically chosen to balance registration accuracy and computational efficiency in C3VD [19]. If a sampled frame lacked informative edges (e.g., textureless regions) or exhibited artifacts (e.g., lens cleaning), a nearby frame with clearer edges was manually selected as a substitute. After selecting keyframes, preprocessing was applied to facilitate the generation of high-quality edge maps. Specular reflections were first removed using Telea inpainting [51] to suppress spurious edges. Due to the diverse range of tissue colors, textures and lighting conditions present in the phantom videos, contrast enhancement was necessary. Keyframes were denoised using a Gaussian filter and converted to grayscale. Subsequently, Contrast Limited Adaptive Histogram Equalization (CLAHE) [52] was applied to enhance local contrast. Then, edge extraction was done using the deep learning-based method DexiNed[53]. The pre-trained model weights (BIPED_10.pt) available from the DexiNed repository were used for this purpose. DexiNed edge maps were inverted (black background, white edges) to match conventional edge formats. Non-maximum suppression and hysteresis thresholding from the Canny edge detector [54] were applied to produce clean binary edge maps. These preprocessing steps generated binary edge maps with thin edges from the input RGB key frames.
@@
The pose log was preprocessed to reduce noise. A box filter with a window size of 54 ms was applied to the data. The pose log was then downsampled to 60 Hz. It is important to note that the pose log records the position of the robotic arm’s end-effector relative to its base, not the camera itself. Hence, a hand-eye calibration was performed to estimate the camera poses. The calibration computed a transformation matrix $\mathbf{X}$ relating the robotic arm motion $\mathbf{A}$ to the colonoscope camera motion $\mathbf{B}$. This relationship was expressed by the equation $\mathbf{A}\mathbf{X}=\mathbf{X}\mathbf{B}$. The Park & Martin [55] optimization method was used to solve for the unknown transformation $\mathbf{X}$. Synchronization between video frames and pose logs was achieved by computing the temporal offset that maximized the correlation between optical flow magnitudes (representing camera motion) and the translation component of the pose log. Finally, after transforming robotic arm poses into camera poses and synchronizing them with the video frames, the camera pose for each keyframe was computed by interpolation between the temporally closest camera poses.
@@
For simulated screening videos, synchronization between video frames and EM tracker poses was achieved by manually identifying temporal correspondence points, fitting a linear regression model (time_sensor = drift × time_video + offset) to correct for clock drift, and interpolating poses to match adjusted frame timestamps. Clock drift arises due to the lack of hardware synchronization between the colonoscope’s video acquisition and the EM tracker, which operate on separate internal clocks. Over time, small differences in their clock frequencies accumulate, leading to misalignment between recorded timestamps. The linear model compensates for this temporal drift, enabling accurate alignment of pose data with corresponding video frames. After correcting timestamps, the camera poses were obtained from the EM tracker poses via hand-eye calibration as described earlier.
@@
An essential feature of clinical colonoscopes is their wide FoV, which is typically around $170^{\circ}$. This feature maximizes the colon surface area visible during a colonoscopy procedure. C3VD incorporated this fisheye effect in its virtual camera model to render the pixel-wise ground truth. As in C3VD, the spherical omnidirectional camera model proposed in Scaramuzza et al. [56] was used in this dataset.
The spherical camera model establishes a one-to-one correspondence between a pixel (u,v)(u",v") on the image plane and a ray direction in 3D world space $(X,Y,Z)$. The Scaramuzza model employs a polynomial projection function to map the ideal undistorted image coordinates $(u,v)$ to the ray direction, by:
@@
where $\lambda$ is a scalar factor, $\alpha_{i}$ are the polynomial coefficients ($\alpha_{1}=0$), and $\rho=\sqrt{u^{2}+v^{2}}$ is the radial distance of the pixel from the image center. To account for lens-sensor misalignment and distortion, the ideal image coordinates are transformed into real distorted image pixels (u,v)(u",v") using the following equation: